# Optimizing ML Concurrent Computation and Communication with GPU DMA Engines

### v1

@@ VI-G Runtime Heuristic for Resource partitioning with ConCCL
For resource partitioning with ConCCL, we recommend a much smaller subset of steps as proposed in Section V-C for baseline C3. Specifically, only the CU-loss slowdown table for a memory-bound GEMM kernel ought to be created. This table helps the runtime identify the optimal number of CUs to remove from concurrent memory-bound kernels for maximum speedups. In our analysis, for MI300X, taking away eight CUs lead to speedups for memory-bound GEMM kernels.
@@ VII-A1 Addressing Memory Interference in C3
Our approach mitigates compute interference by offloading communication to the GPU’s DMA engines. On the MI300X GPU, DMA transfers inherently do not interact with the L1 and L2 caches, effectively eliminating cache interference at these levels. That said, contention for AMD Infinity Cache and HBM bandwidth remains, impacting performance. We leave to future work exploration of techniques such as memory-bandwidth partitioning via memory space (channel) partitioning amongst kernels, memory-aware scheduling [19] to explicitly tackle memory interference. Specifically for latter, C3 taxonomy hints to hardware can aid in memory traffic prioritization.
VII-A2 Extending ConCCL to All-Reduce Collective
@@ VII-B2 Caching Considerations
In realistic applications, usage of DMA offloads can have interesting caching implications. As an example, on MI300X, a DMA transfer will not observe dirty data from a preceding computation which leaves dirty data in L2. In such cases, a cache flush from L2 to AMD Infinity Cache can be necessary before invoking a DMA transfer. Alternately, preceding kernels can proactively write to common cache levels between compute and DMA engines hiding most of this cost.
@@ VII-B3 Inter-Node Communication Considerations
Our study primarily addresses intra-node communication within an MI300X Infinity Platform node. While this is most prominent for inference and small-scale training, large-scale distributed ML training involves inter-node communication as well. In such cases, ML algorithms attempt to use both intra and inter node communication in tandem with attempting to maximize former as much as possible for better performance. In such cases, even with a focus on intra-node communication only, ConCCL can be useful end-to-end.
@@ VIII Related Work
Given the prevalence of C3 in ML, optimized C3 is key to continued scaling of ML. In our work, we methodically study C3, analyze resulting interference and most importantly, without making any invasive changes in existing system software and hardware, we identify techniques to use available software and hardware resources in optimized way to maximize benefit of C3. Computation-Communication Kernel Overlap: Several GPU communication kernel libraries have been developed which improve collective performance in isolation - RCCL [14], MSCCLang [21], and more [22, 23, 24]. However, none of these take any special measures to mitigate any interference where collective kernels execute concurrently with computation kernels which are abundant in ML executions. Our work instead systematically studies execution of RCCL [14] collectives (which incorporates optimizations from other libraries for an AMD MI300X GPU) in concurrence with compute operations and takes necessary measures to mitigate interference when possible. While some prior works study this overlap algorithmically, they do not provide insights into the resource contention therein [11]. In this work, we not only study these contentions, but also develop two optimizations including, schedule prioritization and compute (CU) partitioning, that improve end-to-end performance of C3. While there have been numerous works on improving concurrent kernel execution on GPUs via appropriate resource partitioning [25, 26], prioritization [20] and concurrency-aware kernels [27, 28, 29], they primarily focus on multiple compute kernels. Our work applies these techniques to coarse-grain compute-communication kernels. NanoFlow [3] also performs resource partitioning for C3; however, they create very fine-grained kernels by slicing ML input from micro into nano-batches, and find necessary resource allocation from a huge and complex search space unlike our simple yet useful lookup based heuristic.
Communication Offload: Works which improve C3 performance usually offload communication to dedicated customized accelerators on GPUs, for example, ACE [30] or require extensive hardware support such as compute-capable switches [31]. We consider offload but by leveraging existing data-movers (DMAs), requiring no hardware changes. While dedicated accelerators/compute-enhanced switches can theoretically have lesser memory contention via internal buffering, in practice this is challenging due to buffer overflows which can only be fixed with large buffers that are impractical for large-sized collectives.
@@ Acknowledgment
AMD, the AMD Arrow logo, AMD CDNA, AMD Instinct, AMD ROCm, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.

### v2

@@ VI-G Runtime Heuristic for Resource partitioning with ConCCL
For resource partitioning with ConCCL, we recommend a much smaller subset of steps as proposed in Section V-C for baseline C3. Specifically, only the CU-loss slowdown table for a memory-bound GEMM kernel ought to be created. This table helps the runtime identify the necessary number of CUs to remove from concurrent memory-bound GEMM kernel to maximize speedup. In our analysis, for MI300X, taking away eight CUs lead to speedups for memory-bound GEMM kernels.
@@ VII-A1 Addressing Memory Interference in C3
Our approach mitigates compute interference by offloading communication to the GPU’s DMA engines. On the MI300X GPU, DMA transfers inherently do not interact with the L1 and L2 caches, effectively eliminating cache interference at these levels. That said, contention for HBM bandwidth remains, impacting performance. We leave to future work exploration of techniques such as memory-bandwidth partitioning via memory space (channel) partitioning amongst kernels, memory-aware scheduling [19] to explicitly tackle memory interference. Specifically for latter, C3 taxonomy hints to hardware can aid in memory traffic prioritization.
VII-A2 Accelerating C3 with All-Reduce Collective
@@
Further, we believe that our proposed techniques of schedule prioritization and resource partitioning are also applicable to all-reduce. This is so, as the key insight of the techniques is that prioritizing or providing QoS for kernel with smaller/complimentary resource requirement helps concurrent performance. As all-reduce has this behavior just like all-gather/all-to-all vs. GEMMs (low CU needs, etc.), these techniques can be applicable to all-reduce as well.
@@ VII-B2 Caching Considerations
On MI300X, switching from CU-compute to DMA transfers does not incur any extra cache write-backs. This is so, as L2 is private per XCD (Figure 2), it is already written-back to AMD Infinity Cache at GPU kernel completion. Further, sDMA engines read data from AMD Infinity Cache and write to HBM allowing subsequent GPU kernel to read sDMA transfer’s output. That said, in alternate architectures, preceding GPU kernels can proactively write to common cache levels between compute and DMA engines preventing any cache-related overheads (e.g., cache write-back, flush) from happening on the critical path.
@@ VII-B3 Inter-Node Communication Considerations
Our study primarily addresses intra-node communication within an MI300X Infinity Platform node. While this is most prominent for inference and small-scale training, large-scale distributed ML training involves inter-node communication as well. In such cases, ML algorithms attempt to use both intra and inter node communication in tandem with attempting to maximize former as much as possible for better performance. This is done as intra-node bandwidth is significantly higher than inter-node bandwidth. In such cases, even with a focus on intra-node communication only, ConCCL can be useful end-to-end. Further, large-scale multi-node ML employ hierarchical collectives which break down collectives into intra and inter-node steps[20] and ConCCL can be utilized for intra-node steps.
@@ VII-B5 Power Considerations
VII-B6 Implications of CPU orchestration
As discussed in Section VI-C, DMAs are orchestrated using CPU, the launch/sync cost for transfers are not amortized at smaller sizes. A GPU control-path might help solve this problem and we leave investigating software/hardware optimizations to address this as part of future work.
@@ VIII Related Work
Given the prevalence of C3 in ML, optimized C3 is key to continued scaling of ML. In our work, we methodically study C3, analyze resulting interference and most importantly, without making any invasive changes in existing system software and hardware, we identify techniques to use available software and hardware resources in optimized way to maximize benefit of C3.
@@
Computation-Communication Kernel Overlap: Several GPU communication kernel libraries have been developed which improve collective performance in isolation - RCCL [14], MSCCLang [22], and more [23, 24, 25]. However, none of these take any special measures to mitigate any interference where collective kernels execute concurrently with computation kernels which are abundant in ML executions. While some prior works study this overlap algorithmically, they do not provide insights into the resource contention therein [11]. Other works study C3 scenarios prevalent in DLRM recommendation [26] and Megatron-LM Transformer [27] model executions using both microbenchmarks as well as end-to-end models executions [28]. Specifically, they study all-reduce slowdowns in concurrence with GEMMs, embedding table lookup and other backprop compute operations. Their study finds that all-reduce can slowdown on an average 1.4 in the Transformer model when overlapped with backprop operations, and by up to 6.2 when overlapped with both GEMMs and memory-bound embedding lookup in DLRM. Our work expands this to systematically study execution of collectives (from RCCL [14] which incorporates optimizations from other libraries for an AMD MI300X GPU) in concurrence with compute operations. It studies additional collectives such as all-gather and all-to-all, and furthermore, proposes a taxonomy to incorporate a wide-range of possible C3 scenarios. This taxonomy enables the study of C3 scenarios with different overlap characteristics (e.g., G-long, C-long) and includes communication-computation pairs that are not only prevalent in models today, but can be possible in future models and/or distributed setups. Finally, we not only study these contentions, but also develop two optimizations including, schedule prioritization and compute (CU) partitioning, that improve end-to-end performance of C3 without any additional accelerator and/or hardware changes. While there have been numerous works on improving concurrent kernel execution on GPUs via appropriate resource partitioning [29, 30], prioritization [21] and concurrency-aware kernels [31, 32, 33], they primarily focus on multiple compute kernels. Our work applies these techniques to coarse-grain compute-communication kernels. NanoFlow [3] also performs resource partitioning for C3; however, they create very fine-grained kernels by slicing ML input from micro into nano-batches, and find necessary resource allocation from a huge and complex search space unlike our simple yet useful lookup based heuristic.
Communication Offload: Works which improve C3 performance usually offload communication to dedicated customized accelerators on GPUs. For example, ACE [28] is a custom accelerator for offloading communication which can free up compute units for concurrent compute. It further buffers intermediate, partially reduced data to avoid their reads and writes to memory which reduces memory interference. Similarly, other works require extensive hardware support such as compute-capable switches to offload reduction operations while reducing volume of data moved over network links and memory sub-system for all-reduce [34]. Furthermore, switch offloads still require GPU kernels to orchestrate data movement and in-switch commands, leading to interference with concurrent computation. We consider offload but by leveraging existing data-movers (DMAs), requiring no hardware changes and show C3 benefits with two key ML collectives, all-gather and all-to-all.
@@ Acknowledgment
We thank our colleagues Wenkai Du, Gilbert Lee, Alexander Kaganov, Leo Dong, Anthony Asaro, Padmini Nujetti for many helpful discussions. We also thank Nicholas Malaya, Alex Habeger, Kevin Cheng, Gowri Shankar Guttikonda for help with equipment. Finally, we also thank Gabriel Loh and the anonymous ISPASS reviewers for helping improve the paper. AMD, the AMD Arrow logo, AMD ROCm, AMD Instinct, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.